Use Henderson Hasselback equation:

 

C. Bicarbonate Buffer

3. Prepare 100 mL of 0.200 bicarbonate buffer, pH 10 by calculating the mass of sodium

  bicarbonate and sodium carbonate.

    a. Use the Henderson-Hasselbalch equation to calculate the volume of each stock solution needed. pH = pKa + log [conjugate base] / [weak acid]1

 b. Check your calculations with other students. See the instructor if there is uncertainty.

c. Make the solution and check the pH of a portion of your

buffer solution using the pH meter.

       L = 245μL

  

3. Preparation of Bicarbonate Buffer

 Calculated Volume

CH3COOH:    

KH2PO4:        

 

Actual pH of the buffer: 


1
Expert's answer
2021-09-17T03:36:02-0400

100 ml × 0.2 = 20 ml

Molar mass of sodium carbonate = 105.9888 g/mol

Molar mass of sodium bicarbonate = 84.007 g/mol

= 105.9888 × 84.007/20

= 4.452 × 102

= log 4.452 × 102 = 2.65

pH = 10 + 2.65 = 12.65